# An Ran Chen

### v1

@@ 4.1.2 Collecting Test Coverage
To obtain the test information for fault localization, we need to collect the test execution results, detailed test coverage, and the bug-fix patch. However, there are some limitations in Defects4j. Recent studies (Kabadi et al., 2023; Chen et al., 2023) found that many bugs in Defects4J v1.0 contain tests from the “future” (i.e., added by developers after the bugs were fixed) in their designated buggy commits. This is problematic because the test coverage may contain developers’ knowledge of the bug, which can cause noise and bias in the result of the downstream research. Because of this, we do not directly utilize Defects4J’s data in our study. Instead, since we have the creation date of each bug report, we extract the commit right before the bug report creation from the project’s repository. This commit represents the code that was available to developers when they started to address the bug report. We refer to this commit as a bug report commit to differentiate it from the buggy commit provided by Defects4J.
@@
Due to the above-mentioned issues in Defects4J, our next step is to collect the test results, and detailed test coverage for the bug report commits. To do that, we utilized GZoltar (Campos et al., 2012). GZoltar is a debugging, fault localization, and coverage extraction tool for Java applications that was often applied to analyze Defects4J bugs (Silva et al., 2021; Zhang et al., 2021; Pearson et al., 2017; Küçük et al., 2019). For each crash report, we check out the bug report commit, compile the system, and execute the tests. We utilize the GZoltar CLI tool to implement a script to extract the test coverage and test execution results for the bugs from our study. Since there are many ancient commits, we encountered many challenges in compiling the system and running the tests. For example, since some commits in a project can use different versions of JVM, we had to implement a tool to automatically switch JVM versions when analyzing the commits. There are also many dependency-related issues, where a project uses different versions of a library across commits, but upgrading/downgrading the library version can cause dependency conflicts that lead to compilation errors. In total, we spent over 100 hours compiling and collecting the data. Despite our best efforts, we excluded some bugs where we were not able to execute and collect the test coverage information.
@@
Table 1 provides detailed information on the bugs studied in this work, i.e., bugs from Defect4J 2.0 with valid bug reports. In total, we include 803 bugs from 15 open-source projects. In this work, we target user-reported bugs with stack traces reported, i.e., crash-reporting bugs. Among the 15 projects, there are 89 crash-reporting bugs (i.e. bugs with stack traces in their bug reports), which represents 11.08% of the total set. Of these 89 bugs, we were able to compile and collect the test execution information for 63 bugs. Finally, we excluded three bugs from our study whose root cause is not in a method, as this work focuses on method-level bug localization. The absence of buggy methods can happen in two different situations: (1) The bugfix involves altering lines outside the methods (for example, changing the value of a global variable); or (2) The bugfix includes the creation of new methods, but no existing method is updated. In total, we study 60 bugs. Despite the low percentage of crash reports found in Defects4j, they match with the percentage of bugs with stack traces in other projects (Chen et al., 2021a).
@@ 4.2 Evaluation Metrics
In this paper, we conduct a series of experiments on locating bugs using stack traces. In order to evaluate and compare the results, we utilize a set of well-consolidated metrics in the fault localization field. Previous studies suggest that performing fault-localization at the file level lacks precision (Kochhar et al., 2016) while opting for the statement-level granularity can cause the approach to miss important code context (Parnin and Orso, 2011). Hence, in this paper, we conduct our analysis at the method level.
Mean Average Precision (MAP). MAP is a metric that considers the rank of all the buggy methods in a given project. It is calculated by taking the mean of the Average Precision (AP) across all these bugs, where the following formula defines the AP, and $m$ is the set of buggy files from a given bug report.
$AP=\frac{\sum_{i=1}^{m}{i}/{Pos(i)}}{m}$
Mean Reciprocal Rank (MRR). MRR considers the position where the first buggy file was ranked and takes the mean from all the bug reports. It is computed by the following formula, where K is the set of bug reports and $rank_{i}$ is the position of the first buggy method in the rank.
$MRR=\frac{1}{K}\sum_{i=1}^{K}\frac{1}{rank_{i}}$
Top K. The Top K metric represents the number of bugs in which at least one buggy method is located between the first K best-ranked methods in the approach. According to a previous study (Parnin and Orso, 2011), developers only check a limited amount of suspicious locations. Due to this, we use 1, 3, and 5 as the values for K.
5 Case Study Results
In this section, we discuss the results of our research questions (RQs). For each RQ, we present the motivation, our approach and the results.
@@ 5 Case Study Results
Motivation. Prior research (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011) uses SBFL to assist in debugging, which relies on failing test coverage. The assumption behind SBFL is that the system has at least one failing test covering the bug (i.e. fault-triggering failing test). Intuitively, if fault-triggering tests fail on the bug report commits (the nearest commit when a bug report was reported), we can leverage SBFL techniques to pinpoint the buggy locations. However, it is not clear if fault-triggering tests are available and failing when a bug report is created (Just et al., 2014a; Haben et al., 2023). Therefore, in this RQ, we study whether there are fault-triggering tests (and if these tests fail) at the moment of the bug report creation and their impact on the SBFL’s efficiency.
@@
Approach. To understand whether the test failures are related to the bugs in crash reports, we analyze the number of crash reports that contain failing fault-triggering tests. As discussed in Section 4.1.2, for each bug report, we perform our study on the bug report commits to avoid biases of ”future” tests. In particular, we execute the fault-triggering tests provided by Defects4J and collect the test results with GZoltar. We also investigate whether fault-triggering tests are related to other types of bug reports (without stack traces) and compare the test results with crash reports. A prior study (Chen et al., 2023) identified a set of bugs without stack traces from Defects4J. Therefore, we used their dataset and identified a total of 157 bugs without stack tracks. This data will allow us to draw a comparison between crash report bugs and bugs without stack traces.
@@
To study the impact of missing failing fault-triggering tests on SBFL techniques, we select Ochiai (Abreu et al., 2007) as the baseline since it is widely used and has been shown to perform well on real faults (Pearson et al., 2017; Le et al., 2013; Abreu et al., 2006). We use the detailed coverage and the test results obtained via the Gzoltar execution to apply the Ochiai formula at the method level. The formula calculates the suspiciousness of each code statement, which allows us to rank the methods based on how probable they contain the bug. The SBFL suspiciousness formulas vary for each technique, but all are based on the idea that the code that is more covered by failing tests and less covered by passing tests is more likely to be buggy. The Ochiai formula assigns a suspiciousness score between 0 (not suspect) and 1 (highly suspect) for each piece of code element (e.g., class, method, or statement). If applied at a method level, we calculate the suspicious score of a given method j as:
@@
in which the terms are as defined in Table  2. Each term n in the formula (e.g., $n_{00}$) corresponds to the number of tests that obey specific criteria for the method j for which the suspiciousness score is being calculated. The first criterion determines the method coverage, whether the method is covered by a test ($n_{1x}$) or not ($n_{0x}$). The second criterion indicates the execution status, passing ($n_{x0}$) or failing ( $n_{x1}$). So, if a test covers the method j and its execution fails, for instance, this test will be computed under $n_{11}$. As a calculation example, suppose that we have the following test case results for a given method M1:
@@
  • 2 failing test cases that do not cover M1 ($n_{10}$)
  • 10 passing test cases that cover M1 ($n_{01}$)
@@
Results. Only 3.33% of the crash report bugs contain fault-triggering tests (failing test cases that uncover the bug). Table 3 shows the test execution and fault-triggering test results for the crash report bugs. When comparing with the bugs without stack traces, we obtain the following numbers:
@@
  • Bugs without Stack Traces: 16 out of 157 bugs have fault-triggering tests (10.19%)
@@
Based on the findings, we observed that the majority of the crash report bugs (96.67%) either lack fault-triggering tests or do not trigger them (i.e., fault-triggering tests did not fail). In addition, we can see that the percentage of fault-triggering tests in crash report bugs is significantly less when compared to the set of bugs without stack traces (3.33% versus 10.19%). The reason may be that such exception-related issues are more often triggered during production and not during testing. Another reason may be that developers are more likely to handle exceptions for debugging purposes rather than fault prevention (Shah et al. (2008)). This shift in the use of exception handling as debugging aids increases the likelihood of exception-related issues occurring in production. In such cases, applying traditional SBFL techniques would not be ideal, given the incapacity of the tool to differentiate buggy statements when there are no test failures covering them. Due to the absence of failing fault-triggering tests, the runtime information included in the bug report (i.e., stack traces) serves as the final resource for assistance. In short, studying how to leverage stack traces for fault localization is an important supplement to traditional SBFL techniques that use failing fault-triggering tests.
SBFL techniques may not be effective in addressing crash reports due to the limited presence of fault-triggering failing tests (3.33% of all bug reports).
@@
To study how the absence of fault-triggering tests impacts SBFL, we apply Ochiai to our studied bugs with stack traces. To evaluate the approach, we utilize the metrics presented in Section 4.2, specifically Top-1, Top-3, Top-5, MAP and MRR. We exclude from the Ochiai results the bugs without failing tests, given that the presence of failing tests is necessary for the computation of the suspicioness score. Out of 60 bugs, 23 did not have any failing tests. Note that even if there is a failing test, the test may not be related to the bug that we are interested in (i.e., it is not a fault-triggering test).
@@
On the bug report commit, SBFL performed poorly in all the projects, only being able to locate 2/60 bugs among the Top-5 methods. Table 4 shows the SBFL results applied to the bug report commit for the crash report bugs that contain failing tests. Ochiai could not locate any bug in Top-1 and located only 1 in Top-3 and 2 in Top-5. The highest MAP and MRR values across projects are 0.0693 and 0.0836, respectively, both in the JacksonCore project. Considering that the values of MAP and MRR range between 0 and 1, the obtained values are extremely low. In contrast, a prior study by Chen et al. (2022) that evaluates Ochiai on the original Defects4J benchmark achieved an average MAP of 0.30 and an average MRR of 0.42, which is significantly higher than our results obtained when there is no failing fault-triggering tests.
@@
The MAP and MRR values assess the effectiveness of fault localization techniques in returning relevant results in the top ranking. These metrics are important indicators that demonstrate the usefulness of fault localization techniques. According to a prior survey by Kochhar et al. (2016), 80% of developers consider a fault localization technique successful if it can localize bugs in the top 5 positions. Hence, our finding shows the inefficacy and limitation of SBFL techniques due to the lack of failing tests related to the bug reports.
@@
Due to a lack of failing fault-triggering tests, traditional SBFL techniques have inferior localization results.
@@
Motivation. In the previous RQ, we observed that most crash report bugs do not have fault-triggering tests. Because of that, the stack traces stand out as the best source of execution information. In this RQ, we want to study how stack traces are related to the buggy location and how they can be used to locate these bugs. To do so, we look at two things: (i) what is the type of modification performed in the bug fix to resolve the exception, and (ii) how far away the buggy methods (i.e., methods updated in the bug fix) are from the methods in the stack traces.
@@
Results. In 83% of the cases, the bugfix intention is to prevent the exception from happening again. Figure 7 shows the bugfix intention type distribution for studied bugs. In the majority of the fixes (83.3%), the intention behind the fix is Exception Prevention (i.e., to prevent the exception from happening again). This means that the stack traces are directly related to the root cause of the bug, which suggests that stack trace is an essential source of information to locate the bug. The following two predominant categories are Exception Wrapping (10.0%) and Exception Conversion (5.0%), which also possess a correlation between the fix and the stack traces. In only one bug, COMPRESS-31, the exception-related code was not touched for the fix (Keep Throwing category). After a more detailed investigation, we noticed that, despite this bug being in the Defect4j database, its resolution field was marked as “Not A Problem” on Jira, which explains why nothing was done about the exception.
Figure 7: Bugfix intention type distribution in the studied bugs.
In 78.3% of the bugs, the buggy methods are directly reachable from the stack trace, with an average distance of 0.34 method calls. Upon examining the proximity between stack traces and buggy methods, our analysis reveals that 66.67% of the bugs feature at least one buggy method directly present in the stack trace, indicating a zero distance. Furthermore, a total of 78.3% of the bugs contain buggy methods that are reachable from the stack trace (i.e., distance is zero or more), averaging a very short distance from the stack trace methods of 0.34 method calls. In such scenarios, leveraging stack traces to reconstruct the execution path at the time of the exception can be an effective method for pinpointing the bug. We conduct a manual analysis of the remaining 21.7% of the bugs in which the bugs methods are unreachable from the stack trace. We find that, in these cases, the stack traces only contain external entries (i.e., entries referring to external libraries) or the provided stack traces were incomplete. In short, our findings highlight the potential of using stack traces for fault localization.
In 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace (Exception Prevention, Exception Conversion or Exception Wrapping). In addition, 78.3% of the buggy methods are reachable from the stack traces, having an average distance of 0.34 method calls. This shows that the stack traces are a valuable source of information about the bug.
@@
Motivation. In RQ1, we found that the existing failing tests rarely matched the actual fault-triggering tests. Without fault-triggering tests, the performance of SBFL techniques is greatly affected. In addition, in RQ2, we found that the stack traces are a valuable source of information about the bug. The intention behind the bugfix is usually related to the exception, and the buggy methods are often a short distance away from the stack traces. Prior studies (Chen et al., 2021a, b) show that developers usually rely on stack traces when investigating for bugs, as they provide essential information about the buggy location. Stack traces, similar to fault-triggering tests, carry contextual and execution information associated with the root causes of bugs. In a way, stack traces can be used to represent the coverage of a failing fault-triggering test. Therefore, in this RQ, we aim to investigate how stack traces can be leveraged to complement test cases in locating bugs.
Approach. To understand how stack trace information can contribute to locating buggy methods, we proposed to evaluate an approach called SBEST (Spectrum-Based localization Enhanced by Stack Traces). SBEST is a fault localization approach based on SBFL that incorporates stack trace information with the test coverage data. This technique applies SBFL principles, but instead of using the failing fault-triggering test coverage in the Ochiai formula, it considers the methods that appear on the stack trace entries as the source of the failure. We define a method that is causing a test failure if it appears on any of the frames in a stack trace. Prior studies (Wong et al., 2014a; Chen et al., 2021b) have also shown that the position of methods in the stack trace can be useful in fault localization. Therefore, we incorporate a Stack Trace (ST) score into our approach. We compare our approach with two baselines. We use Ochiai as the first one because it is shown to be one of the best SBFL techniques and performs very well on real faults (Pearson et al., 2017). In addition, we leverage the position of the stack trace as the second baseline. For this, we rank the top entries in the stack trace as more suspicious. Below, we describe our approach in detail.
@@
Spectrum Based (SB) Score. The first one, the SB score, is based on an approach designed to use the information from the stack traces in the fault localization process. It is based on the Ochiai formula, and it utilizes existing test coverage information. However, instead of using the failing tests as the hint of the bug location (since there may not be any test failure or the failure is not related to the reported bug, as found in RQ1), we use the tests covering the methods in the stack trace as the fault-triggering tests. We base this change on the fact that the stack trace represents the system execution at the moment of the failure in the same way that the fault-triggering tests do. More specifically, we utilize the methods present on the first five internal entries from the stack traces to select the failing tests that will be applied to Ochiai’s formula. We base this design on the fact that the higher the stack trace entry, the more probable it is to be related to the bug (Wong et al., 2014b).
The selection of failing tests is determined by counting the number of lines each existing test covers in the Stack Trace methods. We first compute the number of covered lines from these methods for each test by applying the following formula:

### v2

@@ 3.1.2 Collecting Test Coverage
To obtain the test information for fault localization, we need to collect the test execution results, detailed test coverage, and the bug-fix patch. However, there are some limitations in Defects4J. Recent studies (Kabadi et al., 2023; Rafi et al., 2025) found that many bugs in Defects4J v1.0 contain tests from the “future” (i.e., added by developers after the bugs were fixed) in their designated buggy commits. This is problematic because the test coverage may contain developers’ knowledge of the bug, which can cause noise and bias in the result of the downstream research. Because of this, we do not directly utilize Defects4J’s data in our study. Instead, since we have the creation date of each bug report, we extract the commit right before the bug report creation from the project’s repository. This commit represents the code that was available to developers when they started to address the bug report. We refer to this commit as a bug report commit to differentiate it from the buggy commit provided by Defects4J.
@@
Due to the above-mentioned issues in Defects4J, our next step is to collect the test results and detailed test coverage for the bug report commits. To do that, we utilized GZoltar (Campos et al., 2012). GZoltar is a debugging, fault localization, and coverage extraction tool for Java applications that was often applied to analyze Defects4J bugs (Silva et al., 2021; Zhang et al., 2021; Pearson et al., 2017; Küçük et al., 2019). For each crash report, we check out the bug report commit, compile the system, and execute the tests. We utilize the GZoltar CLI tool to implement a script to extract the test coverage and test execution results for the bugs from our study. Since there are many ancient commits, we encountered many challenges in compiling the system and running the tests. For example, since some commits in a project can use different versions of JVM, we had to implement a tool to automatically switch JVM versions when analyzing the commits. There are also many dependency-related issues, where a project uses different versions of a library across commits, but upgrading/downgrading the library version can cause dependency conflicts that lead to compilation errors. In total, we spent over 100 hours compiling and collecting the data. Despite our best efforts, we excluded some bugs where we were not able to execute and collect the test coverage information.
@@
Table 1 provides detailed information on the bugs studied in this work, i.e., bugs from Defect4J 2.0 with valid bug reports. In total, we include 803 bugs from 15 open-source projects. In this work, we target user-reported bugs with stack traces reported, i.e., crash-reporting bugs. Among the 15 projects, there are 89 crash-reporting bugs (i.e., bugs with stack traces in their bug reports), which represent 11.08% of the total set. Of these 89 bugs, we were able to compile and collect the test execution information for 63 bugs. Finally, we excluded three bugs from our study whose root cause is not in a method, as this work focuses on method-level bug localization. The absence of buggy methods can happen in two different situations: (1) The bugfix involves altering lines outside the methods (for example, changing the value of a global variable); or (2) The bugfix includes the creation of new methods, but no existing method is updated. In total, we study 60 bugs. Despite the low percentage of crash reports found in Defects4J, they match the percentage of bugs with stack traces in other projects (Chen et al., 2021a).
@@ 3.2 Evaluation Metrics
Previous studies suggest that performing fault localization at the file level lacks precision (Kochhar et al., 2016), while opting for statement-level granularity can cause the approach to miss important code context (Parnin and Orso, 2011). A survey of developers (Kochhar et al., 2016) further indicates that method-level granularity is the most preferred approach for fault localization tasks. Hence, we conduct our analysis at the method-level in this paper by following prior studies (B. Le et al., 2016; Kang et al., 2024b; Benton et al., 2020; Lou et al., 2021).
Mean Average Precision (MAP). MAP is a popular metric for evaluating the effectiveness of ranking systems, such as fault localization (FL) techniques (Sohn and Yoo, 2017; Chen et al., 2021b; Hirsch and Hofer, 2023). It is the mean of the average precision of all faults. It reflects how well an FL technique ranks faulty (relevant) methods near the top. Precision at a given rank $k$ ($P@k$) is defined as the proportion of faulty methods among the top $k$ methods:
$P@k=\frac{\text{number of faulty methods in top }k\text{ ranks}}{k}$
For a given bug, the Average Precision (AP) is computed as the average of the precision values at each rank where a faulty method appears:
$AP=\frac{1}{M}\sum_{k=1}^{N}P@k\cdot\text{rel}(k)$
where $M$ is the total number of faulty methods, $N$ is the total number of ranked methods, and $\text{rel}(k)$ is 1 if the method at rank $k$ is faulty, and 0 otherwise. Finally, Mean Average Precision (MAP) is obtained by averaging the AP values across all bugs in the dataset:
$MAP=\frac{1}{Q}\sum_{q=1}^{Q}AP_{q}$
where $Q$ is the total number of faults. A higher MAP indicates that faulty methods are consistently ranked higher by the localization technique, reflecting better overall performance.
Mean Reciprocal Rank (MRR). MRR is a ranking metric that evaluates how quickly the first faulty (relevant) method appears in the ranked list. For each fault $q$ (out of $Q$ total faults), let $r_{q}$ denote the rank position of its earliest (highest-ranked) faulty method. The reciprocal rank for that fault is then $1/r_{q}$. Mean Reciprocal Rank (MRR) is calculated as the mean of these reciprocal ranks across all faults:
$\mathrm{MRR}=\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{r_{q}}$
A higher MRR indicates that, on average, faulty methods appear closer to the top of the ranking, reflecting better fault localization performance.
Top K. The Top K metric represents the number of bugs in which at least one buggy method is located between the first K best-ranked methods in the approach. According to a previous study (Parnin and Orso, 2011), developers only check a limited number of suspicious locations. Due to this, we use 1, 3, and 5 as the values for K.
4 Results and Analysis
This section discusses the results of our research questions (RQs). We present the motivation, approach, and results for each RQ.
@@ 4 Results and Analysis
Motivation. Spectrum-Based Fault Localization (SBFL) relies on failing test coverage to identify faulty code, under the assumption that failing tests actually reveal the presence of a bug (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011). However, when a bug report is created, it is often unclear whether the tests that fail at that point are actually related to the bug described in the report-these are known as fault-triggering tests (Just et al., 2014a; Haben et al., 2023; Rafi et al., 2025). This question is especially important for crash-related bugs, where developers may only have the stack trace and available test results to guide debugging. With this in mind, in RQ1, we investigate whether the test failures observed at the bug report commit are actually caused by the bug described in the crash report. We then also explore how the presence or absence of such fault-triggering tests impacts the effectiveness of SBFL in localizing the fault. While analyzing fault-triggering tests that are later added to bug reports would reflect post-localization efforts, our goal in RQ1 is to focus on the reporting stage, characterizing the absence of such tests when the bug is first reported, and suggesting a practical solution for this early debugging scenario.
@@
Approach. To understand whether the test failures are related to the bugs in crash reports, we analyze the number of crash reports that contain failing fault-triggering tests. As discussed in Section 3.1.2, we perform our study on the bug report commits for each bug report to capture the exact state of the project when the crash was first reported and to avoid biases of tests added after the bug was reported. Specifically, we execute all available tests using GZoltar to collect test execution results and coverage data, then compare the coverage of failing tests with the methods mentioned in the crash stack traces and the actual buggy methods identified in the fixes. Building on the prior work by Rafi et al. (2025), which revisited Defects4J to analyze fault-triggering tests and bug timelines, we further inspected the bug reports to determine whether they contain stack traces. Through this analysis, we identified 157 bugs without stack traces, allowing us to compare crash-report bugs with non-crash bugs.
@@
To study the impact of missing fault-triggering tests on SBFL techniques, we select Ochiai (Abreu et al., 2007) as the baseline since it is widely used and has been shown to perform well on real faults (Pearson et al., 2017; Le et al., 2013; Abreu et al., 2006). We utilize the detailed coverage and test results obtained through the Gzoltar execution to apply the Ochiai formula at the method level. The formula calculates the suspiciousness of each code statement, allowing us to rank the methods based on their likelihood of containing the bug. The SBFL suspiciousness formulas vary for each technique, but all are based on the idea that the code that is more covered by failing tests and less covered by passing tests is more likely to be buggy. The Ochiai formula assigns a suspiciousness score between 0 (not suspect) and 1 (highly suspect) for each piece of code element (e.g., class, method, or statement). If applied at a method level, we calculate the suspicious score of a given method j as:
@@
in which the terms are defined in Table  2. Each term n in the formula (e.g., $n_{00}$) corresponds to the number of tests that obey specific criteria for the method j for which the suspiciousness score is being calculated. The first criterion determines the method coverage, whether the method is covered by a test ($n_{1x}$) or not ($n_{0x}$). The second criterion indicates the execution status, passing ($n_{x0}$) or failing ( $n_{x1}$). So, if a test covers the method j and its execution fails, for instance, this test will be computed under $n_{11}$. As a calculation example, suppose that we have the following test case results for a given method M1:
@@
  • 2 failing test cases that do not cover M1 ($n_{01}$)
  • 10 passing test cases that cover M1 ($n_{10}$)
@@
Results. Only 3.33% of the crash report bugs include failing tests that are directly caused by the bug (i.e., fault-triggering tests). Table 3 shows the test execution and fault-triggering test results for the crash report bugs. When comparing with the bugs without stack traces, we obtain the following numbers:
@@
  • Bugs without Stack Traces: 16 out of 157 bugs have fault-triggering tests 10.19%)
@@
Based on the findings, we observed that most crash report bugs (96.67%) either lack fault-triggering tests or do not trigger them (i.e., fault-triggering tests did not fail). In addition, we can see that the percentage of fault-triggering tests in crash report bugs is significantly lower (3.33%) compared to the set of bugs without stack traces (10.19%). The reason may be that such exception-related issues are more often triggered during production, rather than during testing. Another reason may be that developers are more likely to handle exceptions for debugging purposes rather than fault prevention (Shah et al., 2008). This shift in exception handling as a debugging aid increases the likelihood of exception-related issues occurring in production. In such cases, applying traditional SBFL techniques would not be ideal, given the tool’s inability to differentiate between buggy and non-buggy statements when no test failures cover them. Due to the absence of failing fault-triggering tests, the runtime information included in the bug report (i.e., stack traces) is the final resource for assistance. In short, studying how to leverage stack traces for fault localization is an essential supplement to traditional SBFL techniques that use failing fault-triggering tests.
Only 3.33% of crash-report bugs have failing tests that expose the bug (i.e., fault-triggering tests), limiting the applicability of coverage-based fault localization.
@@
To study how the absence of fault-triggering tests impacts SBFL, we apply Ochiai to our studied bugs. To evaluate the approach, we utilize the metrics presented in Section 3.2, specifically Top-1, Top-3, Top-5, MAP, and MRR. We exclude bugs without failing tests from the Ochiai results, given that the presence of failing tests is necessary to compute the suspiciousness score. Out of 60 bugs, 23 did not have any failing tests. Note that even if a test fails, it may not be related to the bug we are interested in (i.e., it is not a fault-triggering test).
@@
On the bug report commit, SBFL performed poorly in all the projects, only being able to locate 2/60 bugs among the Top-5 methods. Table 4 shows the SBFL results applied to the bug report commit for the crash report bugs that contain failing tests. Ochiai could not locate any bug in Top-1 and located only 1 in Top-3 and 2 in Top-5. These two localized bugs are from the JacksonCore and Compress projects, with JacksonCore containing a fault-triggering test. The highest MAP and MRR values across projects are 0.0693 and 0.0836, respectively, in the JacksonCore project. Considering that the values of MAP and MRR range between 0 and 1, the obtained values are extremely low. In contrast, a prior study by Chen et al. (2022) that evaluates Ochiai on the original Defects4J benchmark achieved an average MAP of 0.30 and an average MRR of 0.42, which is significantly higher than our results obtained when there is no failing fault-triggering tests.
@@
The MAP and MRR values assess the effectiveness of fault localization techniques in returning relevant results in the top ranking. These metrics are essential indicators that demonstrate the usefulness of fault localization techniques. According to a prior survey by Kochhar et al. (2016), 80% of developers consider a fault localization technique successful if it can localize bugs in the top 5 positions. Hence, our finding shows the inefficiency and limitations of SBFL techniques due to the lack of failing tests related to the bug reports.
@@
Due to a lack of fault-triggering tests, traditional SBFL techniques have inferior localization results.
@@
Motivation. In the previous RQ, we observed that most crash report bugs lack fault-triggering tests. As a result, stack traces are the only information available to reflect the code execution at the time of failure. In this RQ, we investigate the relationship between the stack traces included in bug reports and the actual buggy code. Specifically, we aim to assess whether stack traces provide meaningful guidance for identifying the faulty location and understanding the nature of the fix. To do this, we analyze two dimensions: (i) the type of modification made in the fix since this reflects how developers interpreted and responded to the exception, and (ii) the proximity between the modified (buggy) methods and the stack trace methods. These dimensions allow us to evaluate the extent to which stack traces directly support fault localization and resolution.
@@
Results. In 83% of cases, the bug fix intention is to prevent the exception from happening again. In the majority of the fixes (83.3%), the intention behind the fix is Exception Prevention (i.e., to prevent the exception from happening again). This means that the stack traces are directly related to the root cause of the bug, suggesting that the stack trace is an essential source of information for locating the bug. The following two predominant categories are Exception Wrapping (10.0%) and Exception Conversion (5.0%), which also possess a correlation between the fix and the stack traces. In only one bug, COMPRESS-31, the exception-related code was not touched for the fix (Keep Throwing category). After a more detailed investigation, we noticed that, despite the bug being in the Defect4j database, its resolution field was marked as ‘Not A Problem‘ on Jira, which explains why no action was taken regarding the exception.
In 78.3% of the bugs, the buggy methods are directly reachable from the stack trace, with an average distance of 0.34 method calls. Upon examining the proximity between stack traces and buggy methods, our analysis reveals that 66.67% of the bugs feature at least one buggy method directly present in the stack trace, indicating a zero distance. Furthermore, a total of 78.3% of the bugs contain buggy methods that are reachable from the stack trace (i.e., distance is zero or more), averaging a very short distance from the stack trace methods of 0.34 method calls. In such scenarios, the short distance between the stack trace and buggy methods greatly reduces the search space for developers or automated tools. This makes stack traces particularly effective for fault localization, as they often point to or near the actual faulty method, even in the absence of explicit failing tests. This observation supports our proposed SBEST approach, which uses stack trace methods to approximate fault-triggering behavior. We conduct a manual analysis of the remaining 21.7% of the bugs for which the buggy methods are unreachable from the stack trace. In these cases, we find that the stack traces only contain external entries (i.e., entries referring to external libraries) or the provided stack traces were incomplete. In summary, our findings underscore the potential of utilizing stack traces for fault localization. These findings indicate that stack-trace proximity is not only common but also practically useful as a substitute for inadequate test coverage in spectrum-based fault localization when such tests are absent.
In 98.3% of the studied bugs, the developer’s fix was applied to one of the methods shown in the exception’s stack trace (Exception Prevention, Exception Conversion, or Exception Wrapping). Additionally, 78.3% of the buggy methods are reachable from the stack traces, with an average distance of 0.34 method calls. This shows that the stack traces are a valuable source of information about the bug.
@@
Motivation. In RQ1, we found that the existing failing tests rarely matched the fault-triggering tests. Without fault-triggering tests, the performance of SBFL techniques is greatly affected. In addition, in RQ2, we found that the stack traces are a valuable source of information about the bug. The intention behind the bugfix is usually related to the exception, and the buggy methods are often a short distance away from the stack traces. Prior studies (Chen et al., 2021a, b) show that developers usually rely on stack traces when investigating bugs, as they provide essential information about the buggy location. Stack traces, similar to fault-triggering tests, carry contextual and execution information associated with the root causes of bugs. In a way, stack traces can represent the coverage of a failing fault-triggering test. Therefore, in this RQ, we aim to investigate how stack traces can be leveraged to complement test cases in locating bugs. We conduct this analysis on the same set of 60 crash-reporting bugs identified earlier.
Approach. To understand how stack trace information can contribute to locating buggy methods, we propose to evaluate an approach called SBEST (Spectrum-Based localization Enhanced by Stack Traces). SBEST is a fault localization approach based on SBFL, incorporating stack trace information with the test coverage data. This technique applies SBFL principles, but instead of using the failing fault-triggering test coverage in the Ochiai formula, it considers the methods that appear on the stack trace entries as the source of the failure. We define a process causing a test failure if it appears on any of the frames in a stack trace. Prior studies (Wong et al., 2014; Chen et al., 2021b) have also shown that the position of methods in the stack trace can be helpful in fault localization. Therefore, our approach incorporates a Stack Trace (ST) score.
We compare our approach with four baselines. The first is Ochiai, a widely adopted and high-performing SBFL technique (Pearson et al., 2017). The second is the stack trace position baseline, where methods appearing higher in the stack trace are ranked as more suspicious. The third is Call Frequency-Based Fault Localization (CFBFL) (Vancsics et al., 2021), which replaces binary hit-based coverage with the frequency of method appearances in unique call stack contexts. CFBFL leverages execution information to capture deeper relationships between stack traces and faulty behavior, making it suitable even when explicit failing tests are unavailable or weak. To fairly incorporate this method into our evaluation, we first examined the program versions that immediately precede the bug report, consistent with our SBEST setup. We then implemented CFBFL using its best-performing variant, JaccardC, as reported in their original study (Vancsics et al., 2021). Lastly, we evaluate the SB_score, a partial version of our approach that incorporates coverage from stack-trace-related methods but excludes stack trace position weighting.
@@
Spectrum Based (SB) Score. The SB score is based on an approach designed to utilize information from stack traces in the fault localization process. It is based on the Ochiai formula and utilizes existing test coverage information. However, instead of using the failing tests as a hint of the bug location (since there may not be any test failure or the failure is not related to the reported bug, as found in RQ1), we use the tests covering the methods in the stack trace as the fault-triggering tests. While we use Ochiai in this study due to its strong performance and wide adoption, our approach is not tied to this specific formula; it is general and can be applied using any SBFL formula.
Our key insight is to use the stack trace as a proxy for failing test behavior by identifying tests that cover the methods appearing in the stack trace. These tests are then treated as fault-triggering inputs, enabling the application of SBFL even in the absence of actual failing tests. While we use the Ochiai formula in our evaluation due to its strong performance and popularity, our approach is not tied to any specific formula. The technique is general and can be applied with any SBFL formula, as the novelty lies in how we construct coverage-like information from stack traces rather than in the formula itself.
We base this change on the fact that the stack trace represents the system’s execution at the moment of failure, in the same way that the fault-triggering tests do. More specifically, we utilize the methods present in the first five internal entries from the stack traces to select the proxy failing tests that will be applied to Ochiai’s formula. We base this design on the fact that the higher the stack trace entry, the more probable it is to be related to the bug (Wong et al., 2014).
The selection of proxy failing tests is determined by counting the number of lines each existing test covers in the Stack Trace methods. We first compute the number of covered lines from these methods for each test by applying the following formula: